OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p) (1 point) p = 6 p = 5 p = 7 p = 12

OpenStudy (anonymous):

-6p + 7 = 6p-9+40-16p 12p-16p+31 = 7 -4p = -24 p = 6

OpenStudy (anonymous):

ty